OpenStudy (anonymous):

Use the Pythagorean Theorem to determine the distance between points A(5, 3) and B(-3, -3).

OpenStudy (anonymous):

i think that u have to use the midpoint formula like i told u the formula and solve it or am i wrong?

OpenStudy (anonymous):

im not sure but I'll try

OpenStudy (anonymous):

ok:)

OpenStudy (amistre64):

subtract one point from the other to get your distances: (5 , 3) -(-3, -3) -------- 8 , 9 ; now the distance is just sqrt(8^2 + 9^2)

OpenStudy (amistre64):

but 3-(-3) = 6 ... not 9 lol

OpenStudy (amistre64):

\[distance = \sqrt{8^2+6^2}\]

OpenStudy (anonymous):

ok i tried but it didn't work im going to put up the answer choices hold on.

OpenStudy (anonymous):

sry amirah but i think u have to use the midpoint formula, amistre am i wrong?

OpenStudy (anonymous):

1) 100 2) 36 3) 64 4) 10

OpenStudy (amistre64):

midpoint does not give you distance; if anything the midpoint will just give you half the distance

OpenStudy (amistre64):

sqrt(8^2 + 6^2) = sqrt(86+36) = sqrt(100) = 10

OpenStudy (amistre64):

..... i really cant type and math at the same time today :/

OpenStudy (anonymous):

i dont know what im saying but if u get half of the distance can u multiply the answer by 2 if im wrong please say because im just wondering if u can do it this way

OpenStudy (amistre64):

if you can determine half the distance then yes, multiply by 2 to get the total; but thats more work than needed :)

OpenStudy (anonymous):

i know but even the sqrt is hard

OpenStudy (amistre64):

sqrt(100) is hard? 10*10 = 100 :)

OpenStudy (anonymous):

ok i figured it was 100 cause of the formula you gave me.. thank you!!!

OpenStudy (anonymous):

i know its easy to work but when u see it its so confusing

OpenStudy (amistre64):

math is a language in its own right; so yes, I understand that it can be confusing

OpenStudy (amistre64):

umm ... the answer is 10 :) the distance between your xs is 8, the distance between your ys is 6 x^2 + y^2 = distance^2 8^2 + 6^2 = d^2 64 + 36 = d^2 100 = d^2 sqrt(100) = sqrt(d^2) 10 = d